OpenStudy (anonymous):

Hekp! you'll get a medal! 17-20 are needed!

OpenStudy (usukidoll):

17-20 users needed?

OpenStudy (anonymous):

hold on! sorry dude!

OpenStudy (anonymous):

\[ y = 2(x-5)(x+2)\] So what do you get if you just foil that?

OpenStudy (anonymous):

i will check! hold on!

OpenStudy (anonymous):

@DarkBlueChocobo

OpenStudy (anonymous):

@Easyaspi

OpenStudy (anonymous):

@Directrix please help man!

OpenStudy (anonymous):

or all.. i have no idea what to do

OpenStudy (anonymous):

Yo, foil that bad boy. Do you know what I mean when I say that?

OpenStudy (anonymous):

yeah! what do you do for the rest? Hey my wifi is about to shut off in 5 minutes, so i gotta go.. is there any way you guys could leave the steps and/or answers here for how to do these? Thank you so much! its a giant help! I've been asking around all day!

OpenStudy (anonymous):

I'll leave a process, sure

OpenStudy (anonymous):

Thanks guys i appreciate it! have an amazing night!

OpenStudy (anonymous):

Since you have the standard form of both equations (the standard form is what you get when you completely factor that equation), \[y = ax^2+bx+c\] you can find the vertex form \[ y=a(x-h)+k\] by rearranging a few terms \[h=-\frac{b}{2a} \ \ ; \ \ k \ \ \ \textrm{is the "leftovers," so to speak}\] http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php The last one is not easily splittable (at least to my eyes) so to find the "intercept form" for, \[y=a(x+p)(x+q)\] you'll have to use the quadratic formula using values from the standard form \[\frac{-b± \sqrt{b^2-4ac}}{2a} \] and the two values ("roots") it gives you will be your p and q (the "±" means you'll subtract that big squareroot for one of the "roots" and add it for the other). (hint: a = 1 for the last one; you will also have one positive root and one negative root. You can tell that because the "c" value is negative. The larger root will be negative as well, since the "b" term is also negative))